# The second data release of the Indian Pulsar Timing Array: Investigating the Effect of Coronal Mass Ejection on PSR J1022+1001 and a Possible Mode Change of PSR J2145-0750

### v1

@@ Abstract
The Indian Pulsar Timing Array (InPTA) has recently published its second data release (DR2), comprising the timing analysis of seven years of data on 27 millisecond pulsars (MSPs), observed simultaneously in the 300-500 MHz (band 3) and 1260-1460 MHz (band 5), using the upgraded Giant Metrewave Radio Telescope (uGMRT). The low-frequency data, particularly in band 3, is highly sensitive to propagation effects such as dispersion measure (DM) fluctuations, which can be imprints of some astrophysical phenomena (scientific outliers). Here, we analyze the two outliers of possible astrophysical origin coming from the band 3 DM time series of two pulsars: PSR J1022+1001, with an ecliptic latitude of -$0.06^{\circ}$, and PSR J2145-0750, one of the brightest MSPs, with multi-component profile morphology. Our study reveals compelling evidence for a coronal mass ejection (CME) event traced in the data of PSR J1022+1001, and reports evidence for a potential mode-changing event in PSR J2145-0750. Extending the analyses presented here to the full sample of InPTA-DR2 pulsars is expected to reveal additional CME events, and possible mode-changing events. Such detections will not only improve our understanding of solar and pulsar magnetospheric plasma interactions but will also enable more accurate modelling of DM variations, leading to improved pulsar timing solutions, which are crucial for high-precision Pulsar Timing Array (PTA) science.
@@ 1 Introduction
High-precision timing of radio pulsars in applications such as the search for low-frequency gravitational waves (GWs) using pulsar timing array experiments (PTAs) often faces the challenge of dealing with shifts in time-of-arrival (ToA) of the radio pulse [1]. Such shifts can be caused by variations in the propagation delay through the ionized inter-stellar medium (IISM) – extreme scattering events [2], scintillations [3, 4], as well as through interplanetary medium (IPM) – coronal mass ejection (CME) [5], solar wind (SW) [6, 7, 8], being a few of the explored phenomenons. Such shifts can also be caused by an intrinsic change in the pulse profile from the usually stable standard profile111which we refer to as template profile. due to glitches [9, 10], magnetospheric origins [11], or mode change [12], but not the stochastic variations. This reveals a systematic shift in the timing residual obtained after a comparison of the observed ToA with that predicted by a model of the pulsar incorporating astrometric, spin, propagation, and binary parameters. Therefore, these “outliers” may indicate astrophysical events of interest, which may be one-off transient events or longer time scale events. A couple of good examples are the profile change event in PSR J1713+0747 reported in 2018 [13], 2020 [14], and 2021 [15] (subsequent interpretations can be found in [16, 11]) and the CME-event reported in PSR J2145-0750 [17].
@@
These profile changes can affect the estimated dispersion measure (DM: a measure of the propagation delay, proportional to the integrated column density of electrons in the inter-stellar and inter-planetary medium along the line-of-sight to the pulsar) of the pulsars. Finding and understanding the causes of such outliers in the DM time series is very important. First, it might lead to serendipitous discovery of interesting astronomical events like profile shape change of the pulsar. Second, it has the potential to improve our understanding of how different astrophysical phenomena impact each other, e.g., coronal mass ejection changing the DM of a pulsar. Such outlier events could also affect noise modelling and degrade the sensitivity to gravitational wave searches [18]. Such studies from one pulsar timing experiment (the Indo-Japanese pulsar timing array (InPTA) [19, 20] in the present case) can potentially alert other PTAs to investigate their data, following the example of the profile change event on PSR J1713+0747. As PTA data can be used not only for the detection of low-frequency GWs, but also be used for various other purposes, like improving solar system ephemeris [21, 22, 23, 24], testing theories of gravity [25, 26, 27, 28], improving the mass measurements of the neutron stars [29] and much more, understanding of outliers in the PTA data is always necessary.
@@
Such outliers can easily be determined through the decade-long monitoring of millisecond pulsars (MSPs) in the PTA experiments. Currently, six such experiments are ongoing with time baselines ranging from 4 years to 28 years (European Pulsar Timing Array (EPTA) [30], the InPTA [19, 20], the Parkes Pulsar Timing Array (PPTA) [31], North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [32], Meerkat Pulsar Timing Array [33] and the Chinese Pulsar Timing Array (CPTA) [34]). The data and resources from these PTA experiments are pooled together by the International Pulsar Timing Array (IPTA) [35] to form the most sensitive data set for the search for GWs. These experiments have reported evidence for the low frequency GWs [36, 37, 38, 33, 39].
@@
Recently, InPTA made its second data release (DR2) public [40] with seven years of data on 27 MSPs. This work investigates outliers in the DM time-series of two pulsars from the InPTA DR2 sample and evaluates their potential astrophysical origin. The two pulsars of interest, highlighted here, are – (a) PSR J1022+1001, which has a very low ecliptic latitude of -0.06, and is known for exhibiting significant solar wind-induced DM variability due to its close proximity to the Sun, making it a natural candidate for probing solar events such as coronal mass ejections; (b) PSR J2145-0750, one of the brightest pulsars in the InPTA sample, shows pronounced short-timescale variations in its DM time series, and its multi-component profile morphology allows sensitive tests for intrinsic mode-changing behaviour.
@@
The outline of this paper is as follows. We provide an overview of the observational strategy of InPTA DR2 and the analysis carried out in this paper, in Section 2. In Section 3 the investigation for the detection of CME using PSR J1022+1001 data is demonstrated, while in Section 4 we describe the analysis for a potential mode change in PSR J2145-0750. The implications of the results and future directions are discussed in Section 5.
@@ 2 Observations and Analysis
The InPTA has been monitoring 27 MSPs routinely using simultaneous observations in the 300-500 MHz (band 3) and 1260-1460 MHz (band 5) with a bandwidth of 200 MHz [41, 40]. The unique subarray capability of the upgraded Giant Metrewave Radio Telescope (uGMRT) [42] is utilized by InPTA to achieve this concurrency. The GMRT antennae are split into two phased arrays with 10 and 15 antennae in band 3 and 5 respectively. The data were recorded with the coherent dedispersion [43] mode at band 3 and using incoherent dedispersion at band 5. Therefore, high precision epoch-wise DM time series is obtained by the InPTA with these data. It may be noted that InPTA is the only PTA experiment which is capable of obtaining instantaneous DM measurements using 300 to 1500 MHz concurrent wide-band observations, unlike other PTA experiments, which utilize multiple epochs to carry out low and high frequency observations. Data on these 27 MSPs over a time baseline of seven years were processed and analyzed. This was published recently as part of InPTA Data release 2 [40]. In this paper, we use the intermediate and final products of this data release for further investigations. The observing setup, observing strategy and data processing used in InPTA are described in [40, 44, 41], where the reader can find additional details.
@@
The starting point of our analysis, in this paper, is the DM time series from InPTA DR2 – the DM for any given pulsar, for any given epoch was evaluated using DMCalc [17]. Instead of treating every DM outlier as a potential candidate, we define outliers of interest using astrophysical criteria tailored to each case study, as will be explained in the following sections. The candidate outlier for PSR J1022+1001 is cross-checked against solar probe data, coronagraph images, and solar wind simulations. Consistency between enhanced electron densities, CME trajectories, and the pulsar’s line of sight is used to validate the association. For PSR J2145-0750, we investigate whether any outlier is instead linked to intrinsic profile evolution. The epoch singled out by the peak-ratio distribution is studied in frequency-resolved folded profiles (300-500 MHz), allowing us to establish if the observed profile departs from the standard template in a manner consistent with mode change. This targeted approach enables us to highlight one astrophysically significant outlier for each of the two pulsars, avoiding the ambiguity of treating all statistical outliers equally. We also briefly examined whether any of the DM outliers could be influenced by terrestrial effects such as lightning or terrestrial gamma-ray flashes (TGFs), which are known to cause short-lived ionospheric perturbations [45] and radio emissions around 300–400 MHz [46], overlapping with our observing band. No consistent correlations were found, suggesting that these effects do not significantly impact our observing bands of interest. In the following sections, we explain the methodology and results in detail for the two pulsars of interest.
@@ 3 Investigating the effect of CMEs on PSR J1022+1001
A CME is a large-scale eruption of magnetized plasma from the Sun’s corona that expels bulk of charged particles into interplanetary space. To assess whether the identified DM outlier in PSR J1022+1001 is associated with such a solar energetic event, we examine its context in a step-wise manner. We first determine the outlying epochs most likely to be affected by CMEs, then identify candidate CME events from solar monitoring databases. Next, we check for corroborating signatures in in situ solar wind measurements, before finally estimating the expected DM excess contributed by such an event. This structured approach allows us to systematically test the solar origin of the observed anomaly.
@@ 3.1 Identifying the possible CME affected observations
We first determined the angular separation between the Sun and the pulsar’s line of sight at all observing epochs. To track solar wind related variability, we subtracted the median DM from the epoch-wise DM values – we will refer to this difference as ‘observational’ DM from here onwards. We then examined the variations in observational DMs as a function of solar elongation. In parallel, we estimated the DM contribution expected from a spherically symmetric solar-wind model using the NE_SWNE\_SW, parameter obtained from the timing solution [40]. This DM is subtracted from its median across the DR2 baseline – we refer to this difference as ‘modelled’ DM. We refer to the difference between the ’modelled’ and ‘observational’ DM as the DM-excess.
[…]
Figure 1: ‘Observational’ DM and ‘modelled’ DM vs separation angle (with respect to Sun) for PSR J1022+1001, with the outlier of interest being marked with red circle. The black points represent ‘observational’ DM, while the blue curve denotes the ‘modelled’ DM from the spherically symmetric solar wind model.
@@ 3.2 CME responsible for the DM outlier
To identify the most plausible CME event(s) that could give rise to such DM-excess, we consulted the DONKI [^2] The Space Weather Database Of Notifications, Knowledge, Information (DONKI), developed at the Community Coordinated Modeling Center (CCMC), is a comprehensive on-line tool for space weather forecasters, scientists, and the general space science community – https://kauai.ccmc.gsfc.nasa.gov/DONKI/. [47] database for solar eruptive events for three days prior to MJD 59800. From this list, we assess which modelled CMEs are likely to intersect the pulsar’s line of sight (LOS), and whether they are simultaneously accessible to any spacecraft capable of providing in situ plasma measurements. Two CME events – cme1, and cme2, emerge as relevant candidates. Both were detected by the Solar TErrestrial RElations Observatory (STEREO)-A spacecraft [48], which at the time was located at 0.95 AU from the Sun with a Heliocentric Earth Ecliptic (HEE) latitude of 0.12and longitude of -21.97. The main properties of these events are summarized in Table 1.
@@
To further evaluate the influence of these CMEs on the pulsar’s LOS, we turn to visualizations of the DONKI simulations performed within the WSA–ENLIL+Cone (WEC) framework [49, 50, 51]. Fig. 2 shows the LOS of PSR J1022+1001 overlaid on the simulation snapshot at 06:00:00 (UTC) on 09 August 2022 (the CME-affected epoch). Given the 6-hour temporal resolution of the simulation, which is much longer than the 17-minute observation duration of the pulsar on that day, this snapshot corresponds to the closest available time to the actual observation time (05:18:39 UTC) on 09 August 2022. The figure displays the positions of the Sun, Earth, satellites, LOS of the PSR J1022+1001, all projected onto the ecliptic plane – the Sun is marked with a white circle at the centre, while the Earth is marked with a yellow circle, 1 AU away from the centre, along the $0^{\circ}$ longitude; the projection of the LOS of PSR J1022+1001 is shown with the yellow line, while those for the satellites and planets are shown with varying colours. The first CME (cme1), associated with a dense, fast outflow, passes beyond both the LOS and STEREO-A, as shown in Fig. 2. The second CME (cme2), however, marginally intersects the LOS during the InPTA observation window, and may represent either a compressed solar wind structure trailing cme1 or a weaker independent event. In either case, it is cme2 that is temporally coincident with the anomalous DM behaviour, because of its influence on the pulsar’s LOS on the observation time. The position of the satellites are obtained from the STEREO Science Center, maintained by the National Aeronautics and Space Administration (NASA) [^3] https://stereo-ssc.nascom.nasa.gov/cgi-bin/make_where_gif.
@@
Figure 2: PSR J1022+1001’s LOS (yellow line) overlaid on the MHD simulation snapshot taken at 06:00:00 (UTC), around the pulsar observation time on 09 August 2022. Locations of various planets and satellites are shown with different symbols as explained in the legends. The dashed spirals are the Parker spirals, the colour code represent the density as per the colorbar. The two CMEs – cme1, and cme2 are also labelled. These snapshots are taken from the MHD simulations in DONKI.
@@
Additional support comes from the Large Angle and Spectrometric Coronagraph – C3 (LASCO/C3 coronagraph) [52] data. The base-difference images generated from the corresponding FITS files are carefully oriented with solar north upwards and the Sun centered in the frame. These reveal large-scale outflows associated with the CME events. For PSR J1022+1001, the CME event of interest, i.e., cme2 manifests as a clear mass outflow aligned toward the pulsar LOS. See Fig. 3. In this figure, the pulsar is marked as a red dot, with a dotted line indicating the direction in which the CME should travel to influence the pulsar’s LOS. The LASCO/C3 field of view extends to $\sim$ 8 from the Sun’s center; PSR J1022+1001 lies at an angular separation of 17, and hence falls outside the image plane. Nevertheless, the inferred trajectory of the CME is consistent with an impact on the LOS.

### v2

# Effects of Coronal Mass Ejection on PSR J1022+1001 and Possible Mode Change of PSR J2145-0750 in the InPTA DR2

@@ Abstract
The Indian Pulsar Timing Array (InPTA) has recently published its second data release (DR2), comprising the timing analysis of seven years of data on 27 millisecond pulsars (MSPs), observed simultaneously in the 300-500 MHz (band 3) and 1260-1460 MHz (band 5), using the upgraded Giant Metrewave Radio Telescope (uGMRT). The low-frequency data, particularly in band 3, is highly sensitive to propagation effects such as dispersion measure (DM) fluctuations, which can be imprints of some astrophysical phenomena (scientific outliers). Here, we analyze the two outliers of possible astrophysical origin coming from the band 3 DM time series of two pulsars: PSR J1022+1001, with an ecliptic latitude of $0.06^{\circ}$, and PSR J2145-0750, one of the brightest MSPs, with multi-component profile morphology. Our study reveals compelling evidence for a coronal mass ejection (CME) event traced in the data of PSR J1022+1001, and reports evidence for a potential mode-changing event in PSR J2145-0750. By contrasting these two cases, we show that DM fluctuations due to CME interacions and intrinsic mode-changing events produce distinct observational signatures, enabling a physically informed classification of scientific outliers in PTA datasets. Extending the analyses presented here to the full sample of InPTA-DR2 pulsars is expected to reveal additional CME events, and possible mode-changing events. Such detections will not only improve our understanding of solar and pulsar magnetospheric plasma interactions but will also enable more accurate modelling of DM variations, leading to improved pulsar timing solutions, which are crucial for high-precision Pulsar Timing Array (PTA) science.
journal: High Energy Astrophysics
@@ 1 Introduction
High-precision timing of radio pulsars in applications such as the search for low-frequency gravitational waves (GWs) using pulsar timing array experiments (PTAs) often faces the challenge of dealing with shifts in time-of-arrival (ToA) of the radio pulse (Verbiest24). Such shifts can be caused by variations in the propagation delay through the ionized inter-stellar medium (IISM) – extreme scattering events (extreme_scat), scintillations (scintillation_lowfreq; scintillation_EPTA), as well as through interplanetary medium (IPM) – coronal mass ejection (CME) (Howard_CME), solar wind (SW) (Caterina_SW_PTA_1; Caterina_SW_PTA_2; Caterina_SW_PTA_3), being a few of the explored phenomenons. Such shifts can also be caused by an intrinsic change in the pulse profile from the usually stable template profile (see Rana25 for the latest template generation scheme), due to glitches (himanshu_glitch_1; himanshu_glitch_2), magnetospheric origins (Rami_profilechange_J1713), or mode change (modechange_J1909_Miles). This reveals a systematic shift in the timing residual obtained after a comparison of the observed ToA with that predicted by a model of the pulsar incorporating astrometric, spin, propagation, and binary parameters. Therefore, these “outliers” may indicate astrophysical events of interest, which may be one-off transient events or longer time scale events. A couple of good examples are the profile change event in PSR J1713+0747 reported in 2018 (Lam_J1713_Profilechange), in 2020 (Goncharov_J1713_ProfileChange), and in 2021 (Singha21) (subsequent interpretations can be found in Jennings_J1713_profilechange; Rami_profilechange_J1713), and the CME-event reported in PSR J2145-0750 (KK_CME_DMCalc).
@@
These astrophysical phenomena can affect the estimated dispersion measure (DM) [^1] It is a measure of the propagation delay, proportional to the integrated column density of electrons in the inter-stellar and inter-planetary medium along the line-of-sight to the pulsar. of the pulsars. Finding and understanding the causes of such outliers in the DM time series is very important. First, it might lead to serendipitous discovery of interesting astronomical events like profile shape change of the pulsar. Second, it has the potential to improve our understanding of how different astrophysical phenomena impact each other, e.g., CME changing the DM of a pulsar. Such outlier events could also affect noise modelling (Aman23) and degrade the sensitivity to gravitational wave searches (Golam_Outliers_GWB). Such studies from one pulsar timing experiment (the Indian pulsar timing array (InPTA) (Joshi18; Joshi22) in the present case) can potentially alert other PTAs to investigate their data, following the example of the profile change event on PSR J1713+0747. As PTA data can be used not only for the detection of low-frequency GWs, but also be used for various other purposes, like improving solar system ephemeris (SolarSystem_PTA_Champion2010; SolarSystem_PTA_Caballero2018; SolarSystemEphemerisError_PTA_Li2016; SolarSystemLINIMOSS_PTA_Guo2019), testing theories of gravity (gravitytest_PTA_1; gravitytest_PTA_2; gravitytest_PTA_3; gravitytest_PTA_4), improving the mass measurements of the neutron stars (neutronstarmass_PTA_1) and much more, understanding of outliers in the PTA data is always necessary.
@@
Such outliers can easily be determined through the decade-long monitoring of millisecond pulsars (MSPs) in the PTA experiments. Currently, six such experiments are ongoing with time baselines ranging from 4 years to 28 years (European Pulsar Timing Array (EPTA) (EPTA), the InPTA (Joshi18; Joshi22), the Parkes Pulsar Timing Array (PPTA) (PPTA), North American Nanohertz Observatory for Gravitational Waves (NANOGrav) (McLaughlin2013), Meerkat Pulsar Timing Array (MPTA) and the Chinese Pulsar Timing Array (CPTA) (CPTA2025)). The data and resources from these PTA experiments are pooled together by the International Pulsar Timing Array (IPTA) (IPTA) to form the most sensitive data set for the search for GWs. These experiments have reported evidence for the low frequency GWs (Agazie2023; epta+inpta2023a; CPTAGW; MPTA; PPTAGW).
@@
Recently, InPTA made its second data release (DR2) public (Rana25) with seven years of data on 27 MSPs. This work investigates outliers in the DM time-series of two pulsars from the InPTA-DR2 sample and evaluates their potential astrophysical origin. The two pulsars of interest, highlighted here, are – (a) PSR J1022+1001, which has a very low ecliptic latitude of -0.06, and is known for exhibiting significant solar wind-induced DM variability due to its close proximity to the Sun, making it a natural candidate for probing solar events such as CMEs; (b) PSR J2145-0750, one of the brightest pulsars in the InPTA sample, shows pronounced short-timescale variations in its DM time series, and its multi-component profile morphology allows sensitive tests for intrinsic mode-changing behavior. Distinguishing the deviation in DM arising from CME interactions, and from intrinsic mode changes, requires that the very definition of an outlier be tailored to the physical scenario under consideration. In this work, we adopt such a scenario-driven approach: the definition of an outlier for PSR J1022+1001 is based on solar elongation geometry and deviations from modelled solar-wind DM, whereas for PSR J2145-0750 it is based on statistically significant changes in its pulse-shape properties. This framework is essential for isolating astrophysically motivated events in PTA datasets, before applying deeper diagnostic analyses to confirm their origin.
@@
The outline of this paper is as follows. We provide an overview of the observational strategy of InPTA-DR2 and the analysis carried out in this paper, in Section 2. In Section 3 the investigation for the detection of CME using PSR J1022+1001 data is demonstrated, while in Section 4 we describe the analysis for a potential mode change in PSR J2145-0750. Section 5 provides a comparative discussion of these two case studies, highlighting the observational signatures that distinguish CME-driven and mode-change-driven DM excursions. The implications of the results and future directions are discussed in Section 6.
@@ 2 Observations and Analysis
The InPTA has been monitoring 27 MSPs routinely using simultaneous observations in the 300-500 MHz (band 3) and 1260-1460 MHz (band 5) bands, as well as single band 3 observations at low frequency, all with a bandwidth of 200 MHz [^2] Band 1 (50-80 MHz) is not yet commissioned by uGMRT, while band 2 (120-250 MHz) is unsuitable for high-precision PTA science due to strong scattering and substantial radio frequency interferences (RFIs). Initial tests showed that band 4 (550–750 MHz or 650-850 MHz) observations in triple band or dual band configuration could not offer any better sensitivity or DM precision, and hence not considered.. The unique subarray capability of the upgraded Giant Metrewave Radio Telescope (uGMRT) (Gupta17) is utilized by InPTA to achieve this concurrency. The uGMRT antennae are split into two phased arrays with 10 and 15 antennae in band 3 and 5 respectively, for band 3-5 dual band observation, while a phased array of 20 antennae are used for the single band 3 observations. The data were recorded with the coherent dedispersion (CDP_Yashwanth) mode at band 3 and using incoherent dedispersion at band 5333Since the low frequency, band 3, is strongly affected by dispersive effects, even a small width of the frequency channels retains small amount of dispersion effect, that remains uncorrected in the incoherent dedispersion used. Hence the real time coherent dedispersion is used to mitigate the limitation due to finite width of the frequency channels in band 3. However, since the high frequency, band 5, is not significantly affected by dispersion, incoherent dedispersion is sufficient to correct for dispersive delays in band 5.. High precision epoch-wise DM time series is obtained by the InPTA with these data. It is notable that InPTA is the only PTA experiment which is capable of obtaining instantaneous DM measurements using 300 to 1500 MHz concurrent wide-band observations, unlike other PTA experiments, which utilize multiple epochs to carry out low and high frequency observations. However, the single band 3 observation data has a relatively higher baseline compared to the dual band 3-5 configuration, and since the lowest frequency band 3 is most sensitive to propagation effects, we work with only band 3 observation data here. Besides, since any intrinsic changes to the pulsar is expected to leave a broad band (in frequency) signature, we should, in principle, be able to capture such intrinsic changes as well in our single band 3 data. In this paper, we use the intermediate and final products of InPTA-DR2 for further investigations. The observing setup, observing strategy and data processing used in InPTA are described in Rana25; Abhimanyu_PINTA; Tarafdar22, where the reader can find additional details.
The starting point of our analysis, in this paper, is the intermediate epoch-wise DM estimates, along with their associated uncertainties [^4] DM uncertainty refers to the confidence interval on our DMCalc-evaluated DM value. In particular, the DM uncertainty is the 1-$\sigma$ formal error returned by a weighted least-squares fit in TEMPO2 (Hobbs_TEMPO2; Edwards_TEMPO2), performed on the cleaned subbanded TOAs – after removal of low-S/N points and Huber-regression-based outliers as implemented in DMCalc., in band 3, produced during the InPTA-DR2 analysis, using DMCalc (KK_CME_DMCalc). Such DM time series contains fluctuations arising from a mixture of propagation effects, and intrinsic pulsar variability. Moreover, because different physical processes imprint qualitatively different signatures on the DM series, defining an outlier purely from a generic statistical deviation is insufficient and may even be misleading. Instead, the astrophysically motivated definition of an outlier is tailored to the physical scenario under investigation, as will be explained in the following sections. In addition to these astrophysical reasons, apparent DM deviations may also arise from observational systematics; however, poor signal-to-noise ratio (S/N) epochs affected by instrumental issues, or other data-quality concerns are systematically excluded through standard post-processing diagnostics and observatory logs, ensuring that only genuine astrophysical candidates are considered. This contextual criterion for defining outliers is crucial: it allows us to isolate physically meaningful events from the DM series before performing the deeper diagnostic analyses that ultimately confirm their astrophysical origin.
@@
The candidate outlier for PSR J1022+1001 is cross-checked against solar probe data, coronagraph images, and solar wind simulations. Consistency between enhanced electron densities, CME trajectories, and the pulsar’s line of sight is used to validate the association. For PSR J2145-0750, the candidate outlier is studied in frequency-resolved folded profiles (300-500 MHz), allowing us to establish if the observed profile departs from the standard template in a manner consistent with mode change. This targeted approach enables us to highlight one astrophysically significant outlier for each of the two pulsars, avoiding the ambiguity of treating all statistical outliers equally. We also briefly examined whether any of these DM outliers could be influenced by terrestrial effects such as lightning or terrestrial gamma-ray flashes (TGFs), which are known to cause short-lived ionospheric perturbations (kuo2015ionospheric) and radio emissions around 300–400 MHz (richard1986results), overlapping with our observing band. No consistent correlations were found, suggesting that these effects do not significantly impact our observing bands of interest. In the following sections, we explain the methodology and results in detail for the two pulsars of interest.
@@ 3 Investigating the effect of CMEs on PSR J1022+1001
A CME is a large-scale eruption of magnetized plasma from the Sun’s corona that expels bulk of charged particles into interplanetary space. To assess whether any observational epoch of PSR J1022+1001 is affected by such a solar energetic event, we examine in a step-wise manner. We first determine, from the DM time series, the outlying epochs most likely to be affected by CMEs, then identify candidate CME events from solar monitoring databases. Next, we check for corroborating signatures in in situ solar wind measurements, before finally estimating the expected DM excess contributed by such an event. This structured approach as detailed in the following subsections, allows us to systematically test the solar origin of the observed anomaly.
@@ 3.1 Identifying the possible CME affected observations
For a low-ecliptic-latitude pulsar such as PSR J1022+1001, where solar-wind variability and CME encounters are expected, a solar-event-induced outlier is naturally defined as a DM enhancement that exceeds the predictions of a solar-wind model within a small solar elongation. This restriction to small elongations is motivated by the fact that, the pulsar signal traverses denser and more variable heliospheric plasma, making any CME-related DM excursion physically identifiable.
Therefore, we first determined the angular separation between the Sun and the pulsar’s line of sight at all observing epochs. We subtracted the median DM from the epoch-wise DM values – we will refer to this difference as ‘observational’ DM from here onwards. We then examined the variations in observational DMs as a function of solar elongation. In parallel, we estimated the DM contribution expected from the widely used spherically symmetric solar-wind model (Caterina_SW_PTA_3) using the NE_SWNE\_SW, parameter obtained from the timing solution of InPTA-DR2 (Rana25). This DM is subtracted from its median across the DR2 baseline – we refer to this difference as ‘modelled’ DM. We refer to the difference between the ‘observational’ and ’modelled’ DM as the DM excess. In Fig. 1 the black points with vertical bars represent the observational DM with its uncertainty, while the modelled DM is represented by the blue curve.
Figure 1: ‘Observational’ DM and ‘modelled’ DM vs separation angle (with respect to Sun) for PSR J1022+1001, with the outlier of interest being marked with red circle. The black points with vertical bars represent ‘observational’ DM with its uncertainty, while the blue curve denotes the ‘modelled’ DM from the spherically symmetric solar wind model.
We flagged as outliers those observational epochs within a conservative $20^{\circ}$ solar elongation threshold for which the observational DM exceeds the modelled DM by more than 3 times the corresponding DM uncertainty. For PSR J1022+1001, this procedure yields a promising candidate outlier – MJD 59800 (09 August 2022), with a separation angle of 17.75and a DM excess of $2.91\times 10^{-3}$ pc/cc (see Fig. 1). The profile for this epoch has reasonably high S/N with good detection – no visible deviation from its template profile – thereby ruling out any instrumental or intrinsic origin for such DM excess.
@@ 3.2 CME responsible for the DM outlier
To identify the most plausible CME event(s) that could give rise to such DM excess, we consulted the DONKI [^5] The Space Weather Database Of Notifications, Knowledge, Information (DONKI), developed at the Community Coordinated Modeling Center (CCMC), is a comprehensive on-line tool for space weather forecasters, scientists, and the general space science community – https://kauai.ccmc.gsfc.nasa.gov/DONKI/. (DONKI_1) database for solar eruptive events for three days prior to MJD 59800. From this list, we assess which modelled CMEs are likely to intersect the pulsar’s line of sight (LOS), and whether they are simultaneously accessible to any spacecraft capable of providing in situ plasma measurements. Two CME events – cme1, and cme2, emerge as relevant candidates. Both were detected by the Solar TErrestrial RElations Observatory (STEREO)-A spacecraft (STEREO_Kaiser_1), which at the time was located at 0.95 AU from the Sun with a Heliocentric Earth Ecliptic (HEE) latitude of 0.12and longitude of -21.97. The main properties of these events are summarized in Table 1.
@@
To further evaluate the influence of these CMEs on the pulsar’s LOS, we turn to visualizations of the DONKI simulations performed within the WSA–ENLIL+Cone (WEC) framework (WSA; ENLIL; WSA_ENLIL_CONE). Fig. 2 shows the LOS of PSR J1022+1001 overlaid on the simulation snapshot at 06:00:00 (UTC) on 09 August 2022 (the CME-affected epoch).
@@
Figure 2: PSR J1022+1001’s LOS (yellow line) overlaid on the MHD simulation snapshot taken at 06:00:00 (UTC), around the pulsar observation time on 09 August 2022. Locations of various planets and satellites are shown with different symbols as explained in the legends. The dashed spirals are the Parker spirals (ParkerSpiral1958), the colour code represent the density as per the colorbar. The two CMEs – cme1, and cme2 are also labelled. These snapshots are taken from the MHD simulations in DONKI.
Given the 6-hour temporal resolution of the simulation, which is much longer than the 17-minute observation duration of the pulsar on that day, this snapshot corresponds to the closest available time to the actual observation time (05:18:39 UTC) on 09 August 2022. The figure displays the positions of the Sun, Earth, satellites, LOS of the PSR J1022+1001, all projected onto the ecliptic plane – the Sun is marked with a white circle at the centre, while the Earth is marked with a yellow circle, 1 AU away from the centre, along the $0^{\circ}$ longitude; the projection of the LOS of PSR J1022+1001 is shown with the yellow line, while those for the satellites and planets are shown with varying colours. The first CME (cme1), associated with a dense, fast outflow, passes beyond both the LOS and STEREO-A, as shown in Fig. 2. The second CME (cme2), however, marginally intersects the LOS during the InPTA observation window, and may represent either a compressed solar wind structure trailing cme1 or a weaker independent event. In either case, it is cme2 that is temporally coincident with the anomalous DM behaviour, because of its influence on the pulsar’s LOS on the observation time. The position of the satellites are obtained from the STEREO Science Center, maintained by the National Aeronautics and Space Administration (NASA) [^6] https://stereo-ssc.nascom.nasa.gov/cgi-bin/make_where_gif.
@@
Additional support comes from the Large Angle and Spectrometric Coronagraph – C3 (LASCO/C3 coronagraph) (LASCO_C3) data. The base-difference images generated from the corresponding FITS files are carefully oriented with solar north upwards and the Sun centered in the frame. These reveal large-scale outflows associated with the CME events. For PSR J1022+1001, the CME event of interest, i.e., cme2 manifests as a clear mass outflow aligned toward the pulsar LOS (see Fig. 3), where the pulsar is marked as a red dot, with a dotted line indicating the direction in which the CME should travel to influence the pulsar’s LOS. The LASCO/C3 field of view extends to $\sim$ 8 from the Sun’s center; PSR J1022+1001 lies at an angular separation of 17, and hence falls outside the image plane. Nevertheless, the inferred trajectory of the CME is consistent with an impact on the LOS.